# TACO: Enhancing Multimodal In-context Learning via Task Mapping-Guided Sequence Configuration

### v1

@@ 7 Conclusion
In this work, we systematically demonstrate the principles and critical importance of task mapping within ICL sequences for enabling effective multimodal ICL in LVLMs. These insights further inspire leveraging task mapping to enhance ICL sequence configuration. To this end, we propose a transformer-based model, TACO, which employs task-aware attention to deeply integrate task mapping into the autoregressive process, thereby optimizing sequence configuration. Experiments show consistent outperformance over SOTA baselines, particularly in generalized-mapping tasks. This study not only presents a practical model but also provides the multimodal ICL community with a new and reliable research direction.
@@
Second, this study does not delve into the role of LVLMs’ internal attention mechanisms and hidden state in capturing and utilizing task mapping. Investigating how task mapping manifests within attention layers could uncover deeper connections between sequence configuration and model reasoning, offering another promising avenue for out future work.
@@ Appendix A Additional Related Works
Configuring ICL sequences. To configure high-quality ICL sequences that bolster multimodal ICL in LVLMs, researchers have explored numerous methods, with metric-centric approaches emerging as the most prominent McIlroy-Young et al. (2024). The most direct metric for both implementation and evaluation is similarity. In this category, methods select ICDs from a demonstration library by comparing their embeddings, an approach widely used in retrieval augmented generation (RAG) systems Luo et al. (2024); Chen et al. (2024b). Retrieval strategies based on semantic entropy have also been applied to tasks demanding more fine-grained selection criteria Wu et al. (2023b); Jeon et al. (2024). To accommodate more complex tasks, several novel metrics have been introduced. Guo et al. (2024) define an influence score, which quantifies the change in model confidence induced by each demonstration, and combine this score with entropy to configure ICL sequences. Bhope et al. (2025) leverages log probabilities of LLM-generated outputs to systematically prune the search space of possible orderings. Although these human-designed metrics can partially capture each ICD’s contribution to latent reasoning, the black-box nature of LVLMs prevents them from fully reflecting the model’s internal inference. For example, similarity-based methods may not provide LLMs with deep task mappings Liu et al. (2021); Li et al. (2024). Approaches designed to mitigate ICD bias can also inadvertently introduce new biases Lyu et al. (2023); Yuan et al. (2024). Model-centric methods have also emerged later, employing multiple models for more demanding selection Wu et al. (2023b); Wang et al. (2024); S. et al. (2024). These methods are not end-to-end and overly focus on ICD selection over ordering. One work closely connected to ours is Yang et al. (2024), which introduces a tiny language model composed of two encoder blocks to automatically select and order ICDs. However, its effectiveness on complex tasks is constrained by a lack of deep insight into task mapping.
@@ Appendix B Formal Theoretical Definition
=$\mathcal{F}\;=\;\bigl{\{}\,f:\mathcal{I}\times\mathcal{Q}\to\mathcal{R}\,\bigr{\}}.$
@@
$p_{\theta}(r\mid S^{n})=\Pr_{M_{\theta}}\bigl{(}r\mid\mathrm{Inst};D_{1},\dots,D_{n};(\hat{I},\hat{Q})\bigr{)},$
where $\;S^{n}=(\mathrm{Inst};D_{1},\dots,D_{n};(\hat{I},\hat{Q}))$ and $D_{i}=(I_{i},Q_{i},R_{i})$.
@@
[…]
@@ C.1 Demonstration Configuring Details
(a) Open-ended VQA: The query $Q_{i}$ is the single question associated with the image $I_{i}$, while the response $R{i}$ is the answer to the question, provided as a short response. For the query sample, $\hat{Q}$ represents the question related to the image $\hat{I}$, and $\hat{R}$ is the expected output of the model.
@@ C.2 In-context Lens
To visualize how LVLMs’ internal token outputs evolve during ICL on specific-mapping tasks, we introduce the in-context lens, an adaptation of the logit lens nostalgebraist (2020). Like its predecessor, in-context lens projects each layer’s final token embedding back into the text vocabulary. Because local mappings in specific-mapping tasks are largely uniform, we can vary task difficulty to distinguish shallow recognition from deep recognition. In the HatefulMemes example, shallow recognition corresponds to the binary classification decision, while deep recognition requires detecting harmful content. We therefore define four anchor categories by selecting representative keywords: "Shallow" represents superficial task understanding, focusing on general or surface-level concepts. Anchor words include "category," "judge," "label," "identify," and "predict." "Deep" indicates a more profound comprehension of the task, capturing nuanced or context-sensitive meanings. Anchor words include "hateful," "offensive," "biased," "harmful," and "inappropriate." "Correct" corresponds to the correct answer for the query sample. "Wrong" represents the incorrect answer, opposite to "Correct." We then compute, for each layer, the relative probability of outputs falling into each category (summing to 100%) and visualize the results as pie charts. Figure 7 shows these charts corresponding to the visualizations in Figure 1(c).
Figure 7: In-context lens visualization results on HatefulMemes.
@@ C.4 Ablation Settings
    • BI on Query Image (BI($\hat{I}$)): Instead of applying blur to the entire ICL sequence, (BI)$\hat{I}$ applies Gaussian blur only to the query image $\hat{I}$. This setting helps isolate the effect of degraded query information on task mapping performance.
@@ C.5.2 Order Sensitivity ($\sigma$)
Then we compute the accuracy for each permuted sequence:
$\text{Acc}\bigl{(}S^{n}_{\text{permute},k}\bigr{)}$ $=\frac{\text{Correct Predictions}}{\text{Total Predictions}},$ (25)
$k$ $=1,2,\ldots,K.$ (26)
@@
Finaly, compute the standard deviation of accuracies as $\sigma$:
@@ C.6 Case Study
Figure 8: Four types of ICL sequences in generalized-mapping task.
@@ D.2 Instruction
The $Inst$ generated by GPT-4o in the main experiment is "You will be provided with a series of image-text pairs as examples and a question. Your task involves two phases: first, analyze the provided image-text pairs to grasp their context and try to deeply think about what the target task is; second, use this understanding, along with a new image and your knowledge, to accurately answer the given question." This content demonstrates great orderliness and can act as a good general semantic guide for ICDs and query sample. This style is named chain-of-thought (CoT).
To incorporate the semantic information of $Inst$ and strengthen task representation during the ICL sequence configuration process, we use GPT-01 to generate simplified versions of these $Inst$ and integrate their embeddings into the task guider, which are indicated by $Inst^{\prime}$. The prompt we use is as follows: "This is an instruction to enable LVLMs to understand and perform a multimodal in-context learning task. Please simplify it by shortening the sentence while preserving its function, core meaning, and structure. The final version should be in its simplest form, where removing any word would change its core meaning". This simplification process allows us to investigate how the semantic information density in the instruction impacts TACO’s sequence configuration ability and the performance of LVLMs in ICL. The results show that simplifying the instruction in a prompt before embedding it in the task guider significantly improves the quality of sequence generation. It also helps to avoid issues caused by too long instructions.
@@
$Inst$ Details
Beginning1 (CoT) You will be provided with a series of image-text pairs as examples and a question. Your task involves two phases: first, analyze the provided image-text pairs to grasp their context and try to deeply think about what the target task is; second, use this understanding, along with a new image and your knowledge, to accurately generate the response of the given query.
Beginning2 (PPT) Construct a dynamic pattern repository from image-text samples, then leverage this framework alongside your knowledge base for concurrent visual analysis and query resolution. The key is parallel processing - your pattern matching and knowledge integration should happen simultaneously rather than sequentially.
@@
Middle (CoT) Now you have seen several examples of image-text pairs. Next, you will be given a question. Your task involves two phases: first, revisit the above image-text pairs and try to deeply think about what the target task is; second, use this understanding, along with a new image and your knowledge, to accurately generate the response of the given question.
End (CoT) Now you have seen several examples of image-text pairs and a question accompanied by a new image. Your task involves two phases: first, revisit the provided examples and try to deeply think about what the target task is; second, use this understanding, the new image and your knowledge to accurately generate the response of the given question.
@@
Table 3: Formats of different instruction types and their corresponding details used in the prompt structure for all VL tasks. (Abbreviated) means that the instruction is a simplified version produced by GPT-o1.
@@ D.3 Prompt Details
Models Prompt details
OpenFlamingo-v2 You will be provided with a series of image-text pairs as examples and a question. Your task involves two phases: first, analyze the provided image-text pairs to grasp their context and try to deeply think about what the target task is; second, use this understanding, along with a new image and your knowledge, to accurately generate the response of the given query. <img><IMG_CONTEXT><|endofchunk|> Query: In what country can you see this? Response: vietnam <img><IMG_CONTEXT><|endofchunk|> Query: Is this a buggy or car? Response: buggy <img><IMG_CONTEXT><|endofchunk|> Query: What is this? Response:
[…]
InternVL2 You will be provided with a series of image-text pairs as examples and a question. Your task involves two phases: first, analyze the provided image-text pairs to grasp their context and try to deeply think about what the target task is; second, use this understanding, along with a new image and your knowledge, to accurately generate the response of the given query. <img><IMG_CONTEXT></img> Query: In what country can you see this? Response: vietnam <img><IMG_CONTEXT></img> Query: Is this a buggy or car? Response: buggy <img><IMG_CONTEXT></img> Query: What is this? Response:
Qwen2VL <|im_start|>system You are a helpful assistant.<|im_end|> <|im_start|>user You will be provided with a series of image-text pairs as examples and a question. Your task involves two phases: first, analyze the provided image-text pairs to grasp their context and try to deeply think about what the target task is; second, use this understanding, along with a new image and your knowledge, to accurately generate the response of the given query. <|vision_start|><|image_pad|><|vision_end|>Query:In what country can you see this? Response: vietnam <|vision_start|><|image_pad|><|vision_end|>Query: Is this a buggy or car? Response: buggy <|vision_start|><|image_pad|><|vision_end|>Query: What is this? Response: <|im_end|> <|im_start|>assistant
Table 4: Prompt details for different models used in the experiments. The table outlines how OpenFlamingov2, IDEFICSv1, InternVL2, and Qwen2VL format their image-text interactions, including examples of image-based questions and short answers. Each model follows a multi-phase task structure, where context is absorbed from previous image-text pairs to answer subsequent questions.
@@
Table 4 provides an overview of the prompt details used for the different models in our experiments. Each model, including OpenFlamingoV2, ICDEFICSv1, InternVL2, and Qwen2VL, employs a structured approach to engage with image-text pairs. The two-phase task requires LVLMs to first absorb information from a series of prompts before utilizing that context to answer subsequent questions related to new images. This method allows for enhanced understanding and reasoning based on prior knowledge and context, which is essential for accurate predictions in VL tasks.
@@ E.1.1 Dataset
For the Open-ended VQA task, we utilize the following datasets: VQAv2, which contains images from the MSCOCO dataset and focuses on traditional question-answering pairs, testing the model’s ability to understand both the image and the question. VizWiz presents a more challenging setting with lower-quality images and questions along with a lot of unanswerable questions, pushing models to handle uncertainty and ambiguity. OK-VQA is distinct in that it requires the model to leverage external knowledge beyond the image content itself to generate correct answers, making it a benchmark for evaluating models’ capacity to integrate outside information.
@@
VL-ICL Bench covers a number of tasks, which includes diverse multimodal ICL capabilities spanning concept binding, reasoning or fine-grained perception. Few-shot ICL is performed by sampling the ICDs from the training split and the query examples from the test split. We choose two image-to-text generation tasks from it, which reflects different key points of ICL. Fast Open MiniImageNet task assigns novel synthetic names (e.g., dax or perpo) to object categories, and LVLMs must learn these associations to name test images based on a few examples instead of their parametric knowledge, emphasizing the importance of rapid learning from ICDs. CLEVR Count Induction asks LVLMs to solve tasks like "How many red objects are there in the scene?" from examples rather than explicit prompts. The ICDs’ images are accompanied by obscure queries formed as attribute-value pairs that identify a specific object type based on four attributes: size, shape, color, or material. Models must perform challenging reasoning to discern the task pattern and generate the correct count of objects that match the query attribute.
The datasets in our experiments are evaluated using task-specific metrics, as summarized in Table 6. For the VQA tasks, Hybrid dataset and VL-ICL bench’s tasks, we use accuracy as the metric to assess the models’ ability to provide correct answers:
$Acc_{a_{i}}=max(1,\frac{3\times{\textstyle\sum_{k\in[0,9]}}match(a_{i},g_{k})}{10}),$ (29)
where $a_{i}$ denotes the model’s generated answer, $g_{k}$ denotes the $k$-th ground true answer. $match(\cdot,\cdot)$ decides whether two answers match; if they match, the result is 1, otherwise it is 0.
@@
Table 6: Evaluation metrics used for each dataset. Accuracy is used for VQA datasets (VQAv2, VizWiz, OK-VQA), self-bulit Hybrid dataset and two VL-ICL Bench’s tasks. CIDEr Vedantam et al. (2015) is used for image captioning datasets (Flickr30k, MSCOCO). ROC-AUC is used for the HatefulMemes classification task.
@@ E.1.2 LVLMs
In recent advances of large vision language models (LVLMs), efficient processing of multimodal inputs, especially images, has become a critical focus. Models like OpenFlamingoV2, IDEFICSv2, InternVL2, Qwen2-VL and GPT-4V implement unique strategies to manage and process visual data alongside textual input.
@@
Qwen2.5-VL reduces the number of visual tokens per image via an MLP-based merger that concatenates and compresses spatially adjacent patch features. A native-resolution ViT first splits an image (e.g. 224 × 224 with patch size 14) into a grid of patch embeddings. Rather than feeding all raw patches into the LLM, Qwen2.5-VL groups each 2 × 2 block of adjacent patch features (four tokens), concatenates them, and projects the result through a two-layer MLP into a single fused token aligned with the LLM’s embedding dimension. This achieves a 4× reduction in sequence length, dynamically compressing image feature sequences of varying lengths.

### v2

@@ 7 Conclusion
In this work, we systematically demonstrate the principles and critical role of task mapping within ICL sequences for enabling effective multimodal ICL in LVLMs. These insights further motivate the use of task mapping to explore more effective ICL sequence configuration strategies that truly align model learning behavior and internal demands. To this end, we propose a transformer-based model, TACO, which employs task-aware attention to deeply integrate task mapping into the autoregressive process, thereby optimizing sequence configuration. Experiments show consistent outperformance over SOTA baselines, particularly in generalized-mapping tasks. This study not only presents a practical model but also provides the multimodal ICL community with a new and reliable research direction.
@@
Our analysis does not examine the internal mechanisms through which task mapping is realized.
This study does not delve into the role of LVLMs’ internal attention mechanisms and hidden state in capturing and utilizing task mapping. Investigating how task mapping manifests within attention layers could uncover deeper connections between sequence configuration and model reasoning, offering another promising avenue for our future work.
Acknowledgements
We extend our sincere gratitude to Prof. Ellie Pavlick from Brown University for her constructive suggestions on the empirical design of this study. Her expertise and insights were invaluable to this research.
@@ Appendix A Additional Related Works
Configuring ICL sequences. To configure high-quality ICL sequences that bolster multimodal ICL in LVLMs, researchers have explored numerous methods, with metric-centric approaches emerging as the most prominent McIlroy-Young et al. (2024). The most direct metric for both implementation and evaluation is similarity. In this category, methods select ICDs from a demonstration library by comparing their embeddings, an approach widely used in retrieval augmented generation (RAG) systems Luo et al. (2024a); Chen et al. (2024b). Retrieval strategies based on semantic entropy have also been applied to tasks demanding more fine-grained selection criteria Wu et al. (2023b); Jeon et al. (2024). To accommodate more complex tasks, several novel metrics have been introduced. Guo et al. (2024) define an influence score, which quantifies the change in model confidence induced by each demonstration, and combine this score with entropy to configure ICL sequences. Bhope et al. (2025) leverages log probabilities of LLM-generated outputs to systematically prune the search space of possible orderings. Although these human-designed metrics can partially capture each ICD’s contribution to latent reasoning, the black-box nature of LVLMs prevents them from fully reflecting the model’s internal inference. For example, similarity-based methods may not provide LLMs with deep task mappings Liu et al. (2021); Li et al. (2024b). Approaches designed to mitigate ICD bias can also inadvertently introduce new biases Lyu et al. (2023); Yuan et al. (2024). Model-centric methods have also emerged later, employing multiple models for more demanding selection Wu et al. (2023b); Wang et al. (2024); Van et al. (2024). These methods are not end-to-end and overly focus on ICD selection over ordering. One work closely connected to ours is Yang et al. (2024), which introduces a tiny language model composed of two encoder blocks to automatically select and order ICDs. However, its effectiveness on complex tasks is constrained by a lack of deep insight into task mapping.
@@ Appendix B Formal Theoretical Definition
$\mathcal{F}\;=\;\bigl\{\,f:\mathcal{I}\times\mathcal{Q}\to\mathcal{R}\,\bigr\}.$
@@
[…]
where $)\;S^{n}=(\mathrm{Inst};D_{1},\dots,D_{n};(\hat{I},\hat{Q}$)) and $D_{i}=(I_{i},Q_{i},R_{i})$.
@@
$\text{Specific‐mapping: }f_{1}=f_{2}=\cdots=f_{n},$
$\text{Generalized‐mapping: }\exists\,i\neq j,\;f_{i}\neq f_{j}.$
@@ C.1 Demonstration Configuring Details
(a) Open-ended VQA: The query $Q_{i}$ is the single question associated with the image $I_{i}$, while the response RiR{i} is the answer to the question, provided as a short response. For the query sample, $\hat{Q}$ represents the question related to the image $\hat{I}$, and $\hat{R}$ is the expected output of the model.
@@ C.2 In-context Lens
To visualize how LVLMs’ internal token outputs evolve during ICL on specific-mapping tasks, we introduce the in-context lens, an adaptation of the logit lens nostalgebraist (2020). Like its predecessor, in-context lens projects each layer’s final token embedding back into the text vocabulary. Because local mappings in specific-mapping tasks are largely uniform, we can vary task difficulty to distinguish shallow recognition from deep recognition. In the HatefulMemes example, shallow recognition corresponds to the binary classification decision, while deep recognition requires detecting harmful content. We therefore define four anchor categories by selecting representative keywords: "Shallow" represents superficial task understanding, focusing on general or surface-level concepts. Anchor words include "category," "judge," "label," "identify," and "predict." "Deep" indicates a more profound comprehension of the task, capturing nuanced or context-sensitive meanings. Anchor words include "hateful," "offensive," "biased," "harmful," and "inappropriate." "Correct" corresponds to the correct answer for the query sample. "Wrong" represents the incorrect answer, opposite to "Correct." We then compute, for each layer, the relative probability of the top three relevant decoded tokens falling into each category (summing to 100%) and visualize the results as pie charts. Figure 7 shows these charts corresponding to the visualizations in Figure 1(c).
Figure 7: Visualization of the evolving internal representations of an LVLM on the HatefulMemes dataset during ICL, analyzed through the in-context lens.
@@ C.4 Ablation Settings
[…]
@@ C.5.2 Order Sensitivity ($\sigma$)
Then we compute the accuracy for each permuted sequence $k=1,2,\ldots,K$:
$\text{Acc}\bigl(S^{n}_{\text{permute},k}\bigr)$ $=\frac{\text{Correct Predictions}}{\text{Total Predictions}}.$ (25)
@@
Finally, compute the standard deviation of accuracies as $\sigma$:
@@ C.6 Case Study
Figure 8: Four types of ICL sequences in the generalized-mapping tasks.
@@ D.2 Instruction
The $Inst$ generated by GPT-4o in the main experiment is "You will be provided with a series of image-text pairs as examples and a question. Your task involves two phases: first, analyze the provided image-text pairs to grasp their context and try to deeply think about what the target task is; second, use this understanding, along with a new image and your knowledge, to accurately answer the given question." This content demonstrates great orderliness and can act as a good general semantic guide for ICDs and the query sample. This style is named chain-of-thought (CoT) Li et al. (2025a).
To incorporate the semantic information of $Inst$ and strengthen task representation during the ICL sequence configuration process, we use GPT-4o to generate simplified versions of these $Inst$ and integrate their embeddings into the task guider, which are indicated by $Inst^{\prime}$. The prompt we use is as follows: "This is an instruction to enable LVLMs to understand and perform a multimodal in-context learning task. Please simplify it by shortening the sentence while preserving its function, core meaning, and structure. The final version should be in its simplest form, where removing any word would change its core meaning." This simplification process allows us to investigate how the semantic information density in the instruction impacts TACO’s sequence configuration ability and the performance of LVLMs in ICL. The results show that simplifying the instruction in a prompt before embedding it in the task guider significantly improves the quality of sequence generation. It also helps to avoid issues caused by too long instructions.
@@
InstInst Details
Beginning1 (CoT) You will be provided with a series of image-text pairs as examples and a question. Your task involves two phases: first, analyze the provided image-text pairs to grasp their context and try to deeply think about what the target task is; second, use this understanding, along with a new image and your knowledge, to accurately generate the response to the given query.
Beginning2 (PPI) Construct a dynamic pattern repository from image-text samples, then leverage this framework alongside your knowledge base for concurrent visual analysis and query resolution. The key is parallel processing - your pattern matching and knowledge integration should happen simultaneously rather than sequentially.
@@
Middle (CoT) Now you have seen several examples of image-text pairs. Next, you will be given a question. Your task involves two phases: first, revisit the above image-text pairs and try to deeply think about what the target task is; second, use this understanding, along with a new image and your knowledge, to accurately generate the response to the given question.
End (CoT) Now you have seen several examples of image-text pairs and a question accompanied by a new image. Your task involves two phases: first, revisit the provided examples and try to deeply think about what the target task is; second, use this understanding, the new image, and your knowledge to accurately generate the response of the given question.
@@
Table 4: Formats of different instruction types and their corresponding details used in the prompt structure for all VL tasks. (Abbreviated) means that the instruction is a simplified version produced by GPT-4o.
@@ D.3 Prompt Details
Each model, including OpenFlamingov2, Idefics2, InternVL2.5, and Qwen2.5VL, employs a structured approach to engage with image-text pairs. The two-phase task requires LVLMs to first absorb information from a series of prompts before utilizing that context to answer subsequent questions related to new images. This method allows for enhanced understanding and reasoning based on prior knowledge and context, which is essential for accurate predictions in VL tasks.
@@ E.1.1 Dataset
For the Open-ended VQA task, we utilize the following datasets: VQAv2, which contains images from the MSCOCO dataset and focuses on traditional question-answering pairs, testing the model’s ability to understand both the image and the question. VizWiz presents a more challenging setting with lower-quality images and questions, along with a lot of unanswerable questions, pushing models to handle uncertainty and ambiguity. OK-VQA is distinct in that it requires the model to leverage external knowledge beyond the image content itself to generate correct answers, making it a benchmark for evaluating models’ capacity to integrate outside information.
@@
VL-ICL Bench covers a number of tasks, which include diverse multimodal ICL capabilities spanning concept binding, reasoning or fine-grained perception. Few-shot ICL is performed by sampling the ICDs from the training split and the query examples from the test split. We choose two image-to-text generation tasks from it, which reflects different key points of ICL. Fast Open MiniImageNet task assigns novel synthetic names (e.g., dax or perpo) to object categories, and LVLMs must learn these associations to name test images based on a few examples instead of their parametric knowledge, emphasizing the importance of rapid learning from ICDs. CLEVR Count Induction asks LVLMs to solve tasks like "How many red objects are there in the scene?" from examples rather than explicit prompts. The ICDs’ images are accompanied by obscure queries formed as attribute-value pairs that identify a specific object type based on four attributes: size, shape, color, or material. Models must perform challenging reasoning to discern the task pattern and generate the correct count of objects that match the query attribute.
The datasets in our experiments are evaluated using task-specific metrics, as summarized in Table 6. For the VQA tasks, Hybrid dataset and tasks in VL-ICL Bench, we use accuracy as the metric to assess the models’ ability to provide correct answers.
@@
Table 6: Evaluation metrics used for each benchmark. Accuracy is used for VQA datasets (VQAv2, VizWiz, OK-VQA), self-built Hybrid dataset, and two tasks in VL-ICL Bench. CIDEr Vedantam et al. (2015) is used for image captioning datasets (Flickr30k, MSCOCO). ROC-AUC is used for the HatefulMemes classification task.
@@ E.1.2 LVLMs
In recent advances of LVLMs, efficient processing of multimodal inputs, especially images, has become a critical focus Luo et al. (2024b); Li et al. (2024a, 2025b); Gu et al. (2025). Models like OpenFlamingov2, Idefics2, InternVL2.5, Qwen2.5VL, and GPT-4V implement unique strategies to manage and process visual data alongside textual input Chen et al. (2025); Pan et al. (2025); Ni et al. (2023); An et al. (2024).
@@
Qwen2.5VL reduces the number of visual tokens per image via an MLP-based merger that concatenates and compresses spatially adjacent patch features. A native-resolution ViT first splits an image (e.g. 224 × 224 with patch size 14) into a grid of patch embeddings. Rather than feeding all raw patches into the LLM, Qwen2.5VL groups each 2 × 2 block of adjacent patch features (four tokens), concatenates them, and projects the result through a two-layer MLP into a single fused token aligned with the LLM’s embedding dimension. This achieves a 4× reduction in sequence length, dynamically compressing image feature sequences of varying lengths.